# RETENTION: Resource-Efficient Tree-Based Ensemble Model Acceleration with Content-Addressable Memory

### v1

@@ IV-A1 Objective and Metric
Since previous works already demonstrate that CAM outperform conventional accelerators, in this work we focus on the CAM-oriented optimization, omitting performance comparison with CPU, GPU, FPGA. To evaluate the effectiveness of RETENTION, we compare the number of required TCAMs against those in existing works. Therefore, we select the reduced number of TCAMs as our evaluation metric, and the higher value represents better performance.
@@ IV-A2 Model Settings
Two ensemble models are selected for case study: Random Forest, representing bagging-based methods, and XGBoost, representing boosting-based methods. Both models are trained on five datasets from the widely used UCI Machine Learning Repository [48], as listed in Table I. Each dataset is divided into training and testing sets in a 7:3 ratio to examine the impact of purity threshold pruning on accuracy. For Random Forest, we utilized the open-source library Ranger [49] for forest construction, setting the number of trees to 100. Purity threshold pruning is applied on Random Forest to reduce model complexity, with tolerance set to {1%, 2%, 3%, 4%, 5%}. On the other hand, we used the official XGBoost library [9] to construct XGBoost with default parameters, including max_depth=6 and num_trees=100.
@@ IV-A4 Simulation
The mapping simulation is conducted by storing root-to-leaf paths into tabular format (i.e., each row is a root-to-leaf path, and every column corresponds to a unique condition), and directly calculate the required number of TCAMs, as no open-source simulator currently supports such functionality. All models are mapped to TCAMs with size $64\times 64$ if not specified.
The subsequent sections present detailed experimental results: Section IV-B analyzes RETENTION’s overall performance. Section IV-C investigates the effects of purity threshold pruning on accuracy and TCAM capacity requirement. Section IV-D evaluates the contributions of ODR and SPC individually. Section IV-E explores the impact of the number of trees. Section IV-F examines the influence of TCAM size. and Section IV-G discusses computational overhead, energy consumption, and throughput.
@@ IV-B Experiment 1: Overall Performance of RETENTION
To evaluate the overall performance of RETENTION, we compare Random Forest models with tolerance set to {1%, 3%, 5%} against baseline methods. As shown in Fig. 5, RETENTION exhibits substantial reductions, achieving 66.25% to 99.96% fewer TCAMs required (equivalent to $2.96\times$ to $2420.81\times$ better space efficiency) in energy-efficient mapping, and 70.83% to 99.70% reduction ($3.43\times$ to $334.48\times$ improvement) in space-efficient mapping. Specifically, setting tolerance to a moderate value of 3% yields $4.35\times$ to $207.12\times$ better space efficiency. These results demonstrate that RETENTION effectively minimizes CAM capacity requirement, offering a promising solution for resource-efficient tree-based model acceleration with CAM.
@@ IV-C Experiment 2: Purity Threshold Pruning
Noticing the extraordinary performance of RETENTION, we break down the framework and evaluate each component separately. This section analyzes the effectiveness of purity threshold pruning by mapping Random Forest models with naive unified mapping. By setting tolerance to {1%, 3%, 5%}, purity threshold pruning alone achieves considerable improvements, reducing CAM capacity requirement by 21.04% to 99.93% ($1.27\times$ to $1357.12\times$ improvement) compared to unpruned models. The relatively lower improvement on the Letter and Wine datasets stems from the increased number of classes, which makes it more challenging for nodes to achieve high purity. Pruning such nodes may result in significant accuracy degradation, which the tolerance mechanism prevents. Consequently, only a few higher-purity nodes can be pruned, thereby limiting the impact of the purity threshold pruning. We further scale tolerance up to {10%, 15%, 20%} to investigate the correlation between tolerance and different datasets. As shown in the left part of Fig. 6, the performance of purity threshold pruning converges at different levels of tolerance, depending on task complexity. Though all can achieve over 90% reduction, simple tasks easily converge with low tolerance ({1%, 3%, 5%}, refer to the red columns), while the convergence for complex tasks occurs with relatively higher tolerance ({10%, 15%, 20%}, refer to the blue columns). However, for imbalanced datasets such as Adult, higher tolerance may lead to a model always favoring the majority class. This should be carefully dealt with by leveraging techniques such as class weights, as discussed in Section III-B. The right part of Fig. 6 exhibits that the user-specified tolerance is highly correlated to testing accuracy loss. Since tolerance is the expected accuracy degradation that is sacrificed to achieve lower model complexity, ideally the testing accuracy loss should be lower than or similar to tolerance. Results show that most of the testing accuracy losses are lower than the predefined tolerance, with only one case slightly higher, suggesting that purity threshold pruning effectively reduces model complexity while ensuring controlled accuracy degradation. Models pruned with higher tolerance also follow this trend, though we omitted the statistics here for clarity.
@@ IV-D Experiment 3: Tree Mapping Scheme
Figure 7: Comparison of different data placement strategies (RF: Random Forest, XGB: XGBoost).
After validating purity threshold pruning, we estimate the effectiveness of the proposed data placement strategies. Fig. 7 presents the results of implementing ODR and SPC on XGBoost and unpruned Random Forest. Since unpruned Random Forest is usually more complex than XGBoost, the improvement of applying ODR and SPC on Random Forest is often more obvious. For energy-efficient mapping, ODR presents more than 50% reduction in all cases. Since #unique_conditions\#unique\_conditions of the models trained on CreditApproval and Letter datasets is an order of magnitude fewer than in other datasets, the improvement observed in the Random Forest trained on these two datasets is relatively limited. For space-efficient mapping, SPC exhibits notable improvement, with over 30% reduction in every case, while FR performs worse than the baseline in half of the cases. The underlying reason is that FR only eliminates completely redundant path segments within a TCAM, whereas significant redundancy is still left behind. Thus, FR beats naive independent mapping only when the redundancy within a single tree is overwhelming. On the other hand, by greedily clustering similar paths together, SPC minimizes redundancy while incurring less runtime computational overhead, consistently outperforming FR and naive independent mapping in all cases. The performance of applying SPC to the Random Forest trained on the CreditApproval dataset is nearly optimal (SPC requires 99 TCAMs, while the theoretical minimum is 98 for 6246 paths). However, the exceptionally strong performance of naive independent mapping on this model limits the potential for further improvement. Experimental results show that implementing the tree mapping scheme alone yields $1.46\times$ to $21.30\times$ better space efficiency, highlighting the effectiveness of both ODR and SPC.
@@ IV-E Experiment 4: Number of Trees
Prior experiments demonstrate that RETENTION is extremely effective in minimizing CAM capacity requirement. In this section, we study the impact of number of trees on RETENTION’s performance. Since XGBoost can achieve higher accuracy with an increased number of trees and appropriate hyperparameter tuning, XGBoost with num_trees set to {100, 300, 500} are evaluated in this experiment. As shown in Fig. 8, the performance of RETENTION tends to improve as num_trees increases because the inefficient issue is amplified by num_trees. However, since the training phase incorporates randomness, and the selected conditions for splitting nodes can strongly affect RETENTION’s performance, the result may not always follow the trend.
@@ IV-F Experiment 5: Different TCAM Sizes
In addition to the number of trees, the size of the TCAM is another factor that may affect RETENTION’s performance. Fig. 9 illustrates the impact of mapping XGBoost to TCAMs with three different sizes. Results suggest that as TCAM size increases, RETENTION yields gradually diminishing improvements in energy-efficient mapping, while offering notable gains in space-efficient mapping. Although the designs of both naive unified mapping and ODR are not directly correlated with TCAM size, it becomes increasingly difficult for ODR to eliminate TCAMs as the size increases, since only those entirely filled with X states can be removed. In some edge cases, such as mapping the models trained on the CreditApproval or Letter datasets to $256\times 256$ TCAMs, ODR may offer no improvement because none of the TCAMs are fully redundant even after reordering. Conversely, increasing the TCAM size can sometimes create partially redundant TCAMs with only a few non-X-state cells. In such cases, ODR can effectively mitigate this issue, allowing these TCAMs to be removed and improving space efficiency. For example, mapping the Letter model to $128\times 128$ TCAMs may benefit from this effect. On the other hand, although improvement increases with TCAM size when RETENTION is applied for space-efficient mapping, CAM utilization actually decreases because larger TCAMs require more X states per row, which leads to greater inefficiency. However, since naive independent mapping assigns each tree to a separate TCAM, larger capacities often fail to be fully exploited, leading to greater performance degradation compared to SPC. Consequently, a smaller TCAM is recommended for RETENTION regardless of the optimization objective, though it must still be large enough to accommodate the longest path in the model.
IV-G Computational Overhead, Energy Consumption, and Throughput
Because prior studies omit the cost of input pre-processing, we analyze the computational overhead, energy consumption, and throughput analytically rather than with hardware/software simulation. For energy-efficient mapping, both naive unified mapping and ODR share the same input sequence among TCAMs in the same column, and a path is stored in a single row to prevent additional computational overhead for result retrieval. For space-efficient mapping, naive independent mapping and SPC require a unique input sequence for each TCAM, leading to higher computational overhead. However, as naive independent mapping stores paths to independent rows, and SPC stores paths in a single TCAM, none of them require further processing for result retrieval. Although FR prevents additional input pre-processing by sharing input sequences among TCAM columns, the paths are no longer stored in the same row after the row elimination process, and the associated computational overhead for result retrieval is proportional to #TCAMs. Therefore, considering computational overhead, the complexity of both ODR and SPC is comparable to previous works. In terms of energy consumption, as RETENTION does not require any additional on-the-fly computation, the reduction in the number of TCAMs is equivalent to energy savings. Furthermore, the energy-saving precharge mechanism proposed in DT2CAM is compatible with ODR, enabling additional reductions in energy consumption. Considering throughput, since RETENTION does not require any hardware modification, and the computational overhead is comparable to or even slightly lower than previous works as the required number of TCAM is dramatically reduced, RETENTION is supposed to offer a similar or even higher throughput than others.
@@ V Conclusion
nvTCAM shows great promise to accelerate tree-based model inference effectively and efficiently, yet the CAM capacity requirement is unacceptably high, and most of the cells are storing the X state for format alignment, which is redundant. In this work, we introduce RETENTION, an end-to-end framework designed to reduce the CAM capacity requirement for tree-based model inference. RETENTION introduces purity threshold pruning to minimize model complexity while ensuring controlled accuracy degradation for bagging-based models. A tree mapping scheme with two data placement strategies, occurrence-based double reordering and similarity-based path clustering, is proposed to further alleviate memory redundancy for energy-efficient mapping and space-efficient mapping. Experimental results show that implementing the tree mapping scheme alone achieves $1.46\times$ to $21.30\times$ better space efficiency, while the full RETENTION framework yields $4.35\times$ to $207.12\times$ improvement with less than 3% accuracy loss. These results demonstrate that RETENTION is extremely effective in reducing CAM capacity requirement, offering a resource-efficient solution for tree-based model acceleration. By implementing RETENTION, resource retention becomes feasible, bringing CAM-based acceleration closer to practicality in resource-constrained environments.

### v2

@@ IV-A1 Objective and Metric
Since previous works [35, 31, 32] already demonstrate that CAM outperforms conventional accelerators, in this work we focus on the CAM-oriented optimization, omitting performance comparison with CPU, GPU, FPGA. To evaluate the effectiveness of RETENTION, we compare the number of required TCAMs against previous works, with the reduced number of TCAMs as our evaluation metric. Therefore, higher values represent better performance.
@@ IV-A2 Model Settings
Two ensemble model types are selected for case study: Random Forest, representing bagging-based models, and XGBoost, representing boosting-based models. Both types are trained on five datasets from the widely used UCI Machine Learning Repository [25], as listed in Table I. Each dataset is divided into training and testing sets in a 7:3 ratio to examine the impact of purity threshold pruning on accuracy. For Random Forest, we utilized the open-source library Ranger [43] for forest construction, setting the number of trees to 100. Purity threshold pruning is applied on Random Forest to reduce model complexity, with tolerance set to {1%, 2%, 3%, 4%, 5%}. On the other hand, we used the official XGBoost library [13] to construct XGBoost with default parameters, including max_depth=6 and num_trees=100.
@@ IV-A4 Simulation
The mapping simulation is conducted by storing root-to-leaf paths into tabular format (i.e., each row is a root-to-leaf path, and every column corresponds to a unique condition), and directly calculate the required number of TCAMs, as no open-source simulator currently supports such functionality. All models are mapped to TCAMs with size $64\times 64$ if not specified. Both energy consumption and latency are evaluated on Intel Xeon Gold 6530 under a single-threaded configuration, using Linux Perf for energy profiling and the C++ std::chrono library for ns-resolution timing. To mitigate system noise, all results are averaged over $10^{9}$ iterations.
The subsequent sections present detailed experimental results: Section IV-B analyzes RETENTION’s overall performance. Section IV-C investigates the effects of purity threshold pruning on accuracy and TCAM capacity requirement. Section IV-D evaluates the contributions of ODR and SPC individually. Section IV-E explores the impact of the number of trees. Section IV-F examines the influence of TCAM size. Finally, Section IV-G discusses RENTENTION’s influence on energy consumption and latency. Detailed experimental statistics of applying RETENTION are summarized in Table III.
@@ IV-B Experiment 1: Overall Performance of RETENTION
To evaluate the overall performance of RETENTION, we compare Random Forest models with tolerance set to {1%, 3%, 5%} against baseline methods. As shown in Fig. 5, RETENTION exhibits substantial reductions, achieving 66.25% to 99.96% fewer TCAMs required (equivalent to $2.96\times$ to $2420.81\times$ capacity reduction) in energy-efficient mapping, and 70.83% to 99.70% reduction ($3.43\times$ to $334.48\times$ improvement) in space-efficient mapping. Specifically, setting tolerance to a moderate value of 3% yields $4.35\times$ to $207.12\times$ improvement. These results demonstrate that RETENTION effectively minimizes CAM capacity requirement, offering a promising solution for resource-efficient acceleration with CAM.
@@ IV-C Experiment 2: Purity Threshold Pruning
Figure 7: Comparison of different data placement strategies (XGB: XGBoost, RF: Random Forest).
Noticing the extraordinary performance of RETENTION, we break down the framework and evaluate each component separately. This section analyzes the effectiveness of purity threshold pruning by mapping Random Forest models with naive unified mapping. By setting tolerance to {1%, 3%, 5%}, purity threshold pruning alone achieves considerable improvement, reducing CAM capacity requirement by 21.04% to 99.93% ($1.27\times$ to $1357.12\times$ improvement) compared to unpruned models. The relatively lower improvement on the Letter and Wine datasets stems from the increased number of classes, which makes it more challenging for nodes to achieve high purity. Pruning such nodes may result in significant accuracy degradation, which the tolerance mechanism prevents. Consequently, only a few higher-purity nodes can be pruned, thereby limiting the impact of purity threshold pruning. We further scale tolerance up to {10%, 15%, 20%} to investigate the correlation between tolerance and different datasets. As shown in the left part of Fig. 6, the performance of purity threshold pruning converges at different levels of tolerance, depending on task complexity. Though all can achieve over 90% reduction, simple tasks easily converge with low tolerance ({1%, 3%, 5%}, refer to the red columns), while the convergence for complex tasks occurs with relatively higher tolerance ({10%, 15%, 20%}, refer to the blue columns). However, for imbalanced datasets such as Adult, higher tolerance may lead to a model always favoring the majority class. This should be carefully dealt with by leveraging techniques such as class weights, as discussed in Section III-B. The right part of Fig. 6 exhibits that the user-specified tolerance is highly correlated to testing accuracy loss. Since tolerance is the maximum expected accuracy degradation that is sacrificed to achieve lower model complexity, ideally the testing accuracy loss should be lower than or similar to tolerance. Results show that most of the testing accuracy losses are lower than the predefined tolerance, with only one case slightly higher, suggesting that purity threshold pruning effectively reduces model complexity while ensuring controlled accuracy degradation. Models pruned with higher tolerance also follow this trend, though we omitted the statistics here for clarity.
@@ IV-D Experiment 3: Tree Mapping Scheme
After validating purity threshold pruning, we estimate the effectiveness of the proposed data placement strategies. Fig. 7 presents the results of implementing ODR and SPC on XGBoost and unpruned Random Forest. Since unpruned Random Forest is usually more complex than XGBoost, the improvement of applying ODR and SPC on Random Forest is often more obvious. For energy-efficient mapping, ODR presents more than 50% reduction in all cases. Since #unique_conditions\#unique\_conditions of the models trained on CreditApproval and Letter datasets is an order of magnitude fewer than in other datasets, the improvement observed in the Random Forests trained on these two datasets is relatively limited. For space-efficient mapping, SPC exhibits notable improvement, with over 30% reduction in every case, while FR performs worse than the baseline in half of the cases. The underlying reason is that FR only eliminates completely redundant path segments within a TCAM, whereas significant redundancy is still left behind. Thus, FR beats naive independent mapping only when the redundancy within a single tree is overwhelming. On the other hand, by greedily clustering similar paths together, SPC minimizes redundancy while incurring less runtime computational overhead, consistently outperforming FR and naive independent mapping in all cases. Although applying SPC to the Random Forest trained on the CreditApproval dataset yields minimum improvement, its performance is actually nearly optimal (SPC requires 99 TCAMs, while the theoretical minimum is 98 for 6246 paths). The restricted improvement stems from the exceptional performance of naive independent mapping on this model, which limits the potential for further improvement. Experimental results show that implementing the tree mapping scheme alone yields $1.46\times$ to $21.30\times$ capacity reduction, highlighting its effectiveness.
@@ IV-E Experiment 4: Number of Trees
Prior experiments demonstrate that RETENTION is extremely effective in minimizing CAM capacity requirement. In this section, we study the impact of number of trees on RETENTION’s performance. Since XGBoost can achieve higher accuracy with an increased number of trees and appropriate hyperparameter tuning, XGBoost with num_trees set to {100, 300, 500} are evaluated in this experiment. As shown in Fig. 8, the performance of RETENTION tends to improve as num_trees increases because the inefficient issue is amplified by num_trees. However, since the training phase incorporates randomness, and the selected conditions for splitting nodes can strongly affect RETENTION’s performance, the results may not strictly follow the trend.
@@ IV-F Experiment 5: Different TCAM Sizes
In addition to the number of trees, the size of the underlying TCAM is another factor that may affect RETENTION’s performance. Fig. 9 illustrates the impact of mapping XGBoost to TCAMs with three different sizes. Results suggest that as TCAM size increases, RETENTION yields gradually diminishing improvements in energy-efficient mapping, while offering notable gains in space-efficient mapping. Although the designs of both naive unified mapping and ODR are not directly correlated with TCAM size, it becomes increasingly difficult for ODR to eliminate TCAMs as the size increases, since only those entirely filled with X states can be removed. In some edge cases, such as mapping the models trained on the CreditApproval or Letter dataset to $256\times 256$ TCAMs, ODR may offer no improvement because none of the TCAMs are fully redundant even after reordering. Conversely, increasing the TCAM size can sometimes create partially redundant TCAMs with only a few non-X-state cells. In such cases, ODR can effectively mitigate this issue, allowing these TCAMs to be removed and lowering capacity requirement. For example, mapping the Letter model to $128\times 128$ TCAMs may benefit from this effect. On the other hand, although improvement increases with TCAM size when RETENTION is applied for space-efficient mapping, CAM utilization actually decreases because larger TCAMs require more X states per row. However, since naive independent mapping assigns each tree to a separate TCAM, larger capacities often fail to be fully exploited, leading to greater performance degradation compared to SPC. Consequently, a smaller TCAM is recommended for RETENTION regardless of the optimization objective, though it must still be large enough to accommodate the longest path in the model when applying SPC.
TABLE III: detailed experimental statistics of RETENTION
Model Type Dataset Tolerance Accuracy Loss Mapping #TCAMs (Mapping/Baseline) Reduction Required Capacity (KB)
ODR 7156 / 1482120 $207.12\times$ 3578.0
Random Forest Adult 3% 2.86% FR 2313 / 41613 $7.99\times 11$56.5
SPC 495 / 41613 $84.07\times$ 247.5
ODR 33 / 3038 $92.06\times$ 16.5
Random Forest CreditApproval 3% 1.14% FR 26 / 144 $5.54\times$ 13.0
SPC 12 / 144 $12.00\times$ 6.0
ODR 2833 / 537419 $189.70\times$ 1416.5
Random Forest DryBean 3% 2.76% FR 824 / 7693 $9.34\times$ 412.0
SPC 113 / 7693 $68.08\times$ 56.5
ODR 4783 / 20825 $4.35\times$ 2391.5
Random Forest Letter 3% 2.26% FR 4278 / 15120 $3.53\times$ 2139.0
SPC 1641 / 15120 $9.21\times$ 820.5
ODR 28322 / 144824 $5.11\times$ 14161.0
Random Forest Wine 3% 0.23% FR 8424 / 21608 $2.57\times$ 4212.0
SPC 1163 / 21608 $18.58\times$ 581.5
ODR 227 / 486 $2.14\times$ 113.5
XGBoost Adult - 0.00% FR 162 / 101 Failed 81.0
SPC 54 / 101 $1.87\times$ 27.0
ODR 37 / 85 $2.30\times$ 18.5
XGBoost CreditApproval - 0.00% FR 32 / 100 $3.13\times$ 16.0
SPC 17 / 100 $5.88\times$ 8.5
ODR 3382 / 7790 $2.30\times$ 1691.0
XGBoost DryBean - 0.00% FR 908 / 654 Failed 454.0
SPC 206 / 654 $3.17\times$ 103.0
ODR 1321 / 2960 $2.24\times$ 660.5
XGBoost Letter - 0.00% FR 1292 / 2600 $2.06\times$ 646.0
SPC 740 / 2600 $3.51\times$ 370.0
ODR 2722 / 5565 $2.04\times$ 1361.0
XGBoost Wine - 0.00% FR 1117 / 711 Failed 558.5
SPC 267 / 711 $2.66\times$ 133.5
TABLE IV: energy consumption and latency
Operation (Per TCAM) Energy Latency
Query Packing (Random) $6.27\mu J$ $14.99ns$
Query Packing (Sequential) $5.34\mu J$ $14.70ns$
TCAM Search [12] $2.09pJ$ $0.96ns$
IV-G Experiment 6: Energy Consumption and Latency
The workflow of CAM-based inference is (1) encode input features into a binary sequence, (2) pack and distribute queries to CAMs, (3) search in CAMs, and (4) generate a prediction based on the search results. Since (1) and (4) are mapping-independent, here we only discuss the energy consumption and latency for (2) and (3). Table IV shows the detailed statistics for query packing and TCAM search with size $64\times 64$, measured with the largest #unique_conditions\#unique\_conditions (i.e., 5848) observed from the models listed on Table III. As the thresholds for each feature are sorted for fast encoding, mapping with clustering or column reordering (i.e., ODR, SPC, naive independent mapping) requires random accessing the encoded sequence during query packing, while naive unified mapping accesses sequentially. However, even the longest encoded sequence occupies only 5.71KB, and therefore it can be perfectly cached with minimal influence on latency.
As discussed in Section III-D, naive unified mapping and ODR both require $\left\lceil\frac{\#\textit{unique\_conditions}}{S}\right\rceil$ times of query packing given the size of CAM $S\times S$. Therefore, ODR substantially reduces capacity requirement at the cost of slightly higher energy consumption and latency than naive unified mapping, as the savings from reducing search operations is negligible compared to the cost of random accessing. On the other hand, since both SPC and naive independent mapping require #TCAM times of query packing with random accessing, SPC is more efficient than naive independent mapping considering capacity requirement, energy consumption, and latency.
@@ V Conclusion
nvTCAM shows great promise to accelerate tree-based model inference effectively and efficiently, yet the capacity requirement of CAM is unacceptably high, and most of the cells are storing the X state for format alignment, which is redundant. In this work, we introduce RETENTION, an end-to-end framework designed to reduce CAM capacity requirement for tree-based model inference. RETENTION introduces purity threshold pruning to minimize model complexity while ensuring controlled accuracy degradation for bagging-based models. A tree mapping scheme with two data placement strategies, occurrence-based double reordering and similarity-based path clustering, is proposed to further alleviate memory redundancy for energy-efficient mapping and space-efficient mapping. Experimental results show that implementing the tree mapping scheme alone achieves $1.46\times$ to $21.30\times$ capacity reduction, while the full RETENTION framework yields $4.35\times$ to $207.12\times$ improvement with less than 3% accuracy loss. These results demonstrate that RETENTION is extremely effective in reducing CAM capacity requirement, offering a resource-efficient solution for tree-based model acceleration. By implementing RETENTION, resource retention becomes feasible, bringing CAM-based acceleration closer to practicality in resource-constrained environments.